OpenStudy (anonymous):

you roll 2 dice. What is the probability that the total of the two dice is even or that the total is less than 5?

OpenStudy (amistre64):

less than 5 1 2 3 4 5 6 1 x x x 2 x x 3 x 4 5 6 include all the even sums: 1 2 3 4 5 6 1 x x x x 2 x x x x 3 x x x 4 x x x 5 x x x 6 x x x

OpenStudy (amistre64):

20/36 = 5/9 perhaps?

OpenStudy (amistre64):

if you do the formula, you have to subtract the ones that you count double. those P(A and B) stuff

OpenStudy (amistre64):

P(A or B) = P(A) + P(B) - P(A and B) P(x<5; or x=even) = 6/36 + 18/36 - 4/36 = 2/36 + 18/36 = 20/36 = 5/9

OpenStudy (anonymous):

Thanks!